# Neural Diffusion Models

### v2

@@ 1 Introduction
Generative models are a powerful class of probabilistic machine learning models with a wide range of applications from e.g. art and music to medicine and physics (Tomczak, 2022; Creswell et al., 2018; Papamakarios et al., 2021; Yang et al., 2022). Generative models learn to mimic the underlying probability distribution of a given data set and can generate novel samples that are similar to the original data. They can for example be used for data augmentation, generating synthetic data sets that increase diversity and scale of the training data, as well as for unsupervised learning, discovering patterns and latent structures in data.
Diffusion models have emerged as a family of generative models that excel at several generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020). They parameterize the data model through an iterative refinement process, the reverse process, that builds up the data step-by-step from pure noise. For training purposes an auxiliary noising process, the forward process, is introduced that successively adds noise to data. The reverse process is then optimized to resemble the forward process. Despite success in various domains (Sohl-Dickstein et al., 2015; Ho et al., 2020; Saharia et al., 2021; Popov et al., 2021; Watson et al., 2022; Trippe et al., 2023), a key limitation of most existing diffusion models is that they rely on a fixed and pre-specified forward process that is unable to adapt to the specific task or data at hand. At the same time there are many works (Hoogeboom & Salimans, 2022; Rombach et al., 2022; Lipman et al., 2022) that improve performance of diffusion models by modifications of the forward processes.
@@
The main motivation for NDMs is improved likelihood and density estimation, crucial for applications to compression (MacKay, 2003), semi-supervised learning (Dai et al., 2017), adversarial purification (Song et al., 2018), and many others. However, for completeness we also study the impact of NDMs on image generation quality. We find that for small to medium number of steps NDMs achieves better image generation quality than denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020), and comparable results for a large number of steps.
@@
DDPM (Ho et al., 2020) / DDIM (Song et al., 2020a) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$
Flow Matching OT (Lipman et al., 2022) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$ $\alpha_{t}=t$, $\sigma_{t}=1-(1-\sigma_{\mathrm{min}})t$
@@ 2 Background
The combination of the forward process $q$ and the reverse process $p_{\theta}$ is a form of (hierarchical) variational autoencoder (Kingma & Welling, 2013; Rezende et al., 2014). Therefore, it can be trained by optimizing the usual variational bound on the negative log-likelihood. In the case of diffusion models, it can be written as follows (see Section A of Ho et al. (2020)):
@@
By choosing a specific value for $\tilde{\sigma}_{s|t}^{2}$, we can obtain equality between the processes of DDPM and DDIM (see section 4.1 of Song et al. (2020a)). Furthermore, as Song et al. (2020b) demonstrated, when the number of steps $T$ in DDPM goes to infinity, we can transition to continuous time. In this scenario, the reverse process can be described using a Stochastic Differential Equation (SDE):
@@ 3 Neural diffusion models
Diffusion models can be viewed as a special type of hierarchical variational autoencoders, where the latent variables are inferred using scaling of data points and injecting of Gaussian noise. However, this formulation limits diffusion models in terms of the flexibility of the latent space, which prevents from learning more useful distributions for the reverse (generative) process. To overcome this limitation, we propose a general form of transformations of data that allows to define and learn distributions of the latent space.
@@ 3.1 Model definition and variational objective
for $0\leq s\leq t\leq T$ where $\tilde{\sigma}_{s|t}^{2}\leq\sigma_{s}^{2}$ is a design choice. Using this posterior we can define an implicit forward process according to (1) (see Figure 0(b)). This forward process provides access to both marginal and posterior distributions just like in the DDIM framework (Song et al., 2020a). The corresponding NDM variational objective has the following form:
@@ 3.2 Continuous time NDMs
Similar to training a discrete time NDM, we can train a continuous time NDMs by sampling time. In our experiments we use importance sampling (Song et al., 2021) and sample time from a distribution proportional to $\frac{1}{g^{2}(t)}$.
@@ 4 Experiments
Table 2: Summary of our findings for density modeling tasks.
@@
Score SDE (Song et al., 2020b) $2.99$
Score Flow (Song et al., 2021) $2.83$ $3.76$
@@ 4.1 Implementation details
We demonstrate NDMs with learnable transformations on the MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), downsampled ImageNet (Deng et al., 2009) and CelebA-HQ-256 (Karras et al., 2017) datasets. In all experiments we use same neural network architectures to parameterize both the generative process and the transformations $F_{\varphi}$. In experiments with images we use the U-Net architecture from Dhariwal & Nichol (2021). To ensure consistency with Song et al. (2020b, 2021), we apply horizontal flipping as a data augmentation technique for training models on CIFAR-10 and ImageNet. Unless otherwise stated, we utilize the DDPM variance-preserving schedule of noise injection for $\alpha_{t}$ and $\sigma^{2}_{t}$. For density estimation of discrete data we use uniform dequantization.
In the experiments we report negative log-likelihood (NLL) in bits per dimension (BPD), negative evidence lower bound (NELBO) (8), and sample quality as measured by the Frechet Inception Distance (FID) (Heusel et al., 2017). We calculate NLL by integrating the corresponding ODEs using the RK45 solver from Dormand & Prince (1980), and both NLL and NELBO are calculated on test data. For FID we report the score averaged over 50k images.
@@ 4.2 Learned transformations
Let us examine some of the transformations that NDM learns. Figure 1(a)-1(c) illustrates the transformations that NDM learns for the 2D checkerboard distribution, MNIST, and CIFAR-10 datasets. For the checkerboard, we observe that $F_{\varphi}$ learns to transform the interleaved pattern into a non-interleaved one. In the case of the grayscale digits of the MNIST dataset, $F_{\varphi}$ learns to highlight the distinctive features of the numbers. It thickens the lines and even creates bubbles at the corners. For the color images of CIFAR-10, $F_{\varphi}$ learns to increase the image contrast. In all cases, our model learns a way to simplify the data distribution. These transformations may enable the reverse process to transition more smoothly from simple distributions to complex ones.
@@ 4.3 Image generation
To the best of our knowledge, the NLL results for ImageNet and CelebA-HQ constitutes new state-of-the-art results for diffusion models.
@@
See Appendix B.5 for further discussion and Appendix D for additional results and ablation studies.
@@ 5 Related work
NDMs build on diffusion probabilistic models originally proposed by Sohl-Dickstein et al. (2015), which can be considered as an instance of (hierarchical) variational autoencoders (VAEs) (Kingma & Welling, 2013; Rezende et al., 2014). Recently, the theory of diffusion models was extended to deterministic sampling (Song et al., 2020a) and continuous time (Song et al., 2020b). These results allowed to reach impressive performance in image generation tasks (Ho et al., 2020; Song et al., 2020b; Dhariwal & Nichol, 2021; Kingma et al., 2021). However, most existing diffusion models have a significant limitation in that they rely on a pre-specified and simple noise injection process that is unable to adapt to the specific task or data at hand. To overcome this, researchers have explored ways to generalize diffusion models.
Various papers have since proposed ways to speed up sampling from diffusion models. Tachibana et al. (2021) and Liu et al. (2022a) proposed alternative SDE and ODE solvers. Xiao et al. (2021) proposed replacing simple Gaussian distributions at each generation step with distributions learned by GANs (Goodfellow et al., 2014). Some works proposed methods like iterative distillation with a reduction in the number of steps (Salimans & Ho, 2022) and iterative straightening of trajectories (Liu et al., 2022b; Liu, 2022). While these methods change the generative process, they are compatible with NDMs.
Several papers proposed constructing the process of data corruption not by noise injection, but rather by blurring (Rissanen et al., 2022; Daras et al., 2022; Hoogeboom & Salimans, 2022) or through another linear transformation (Singhal et al., 2023). Another line of work modifies directly the dynamics of diffusion models through mapping the data into the latent space of VAE (Vahdat et al., 2021; Rombach et al., 2022), hierarchical VAE (Gu et al., 2022) or normalizing flow (Kim et al., 2022) models and then runs standard linear diffusion. As we demonstrate in Tables 1, these arise as distinct special cases of NDMs for specific choices of the transformation $F_{\varphi}$.
@@
Inspired by diffusion models, several works (Lipman et al., 2022; Neklyudov et al., 2022) have proposed simulation-free objectives for training continuous normalizing flows. These approaches are similar to diffusion models as they rely on the idea of reversing a predefined corruption process. Later, some works (Albergo & Vanden-Eijnden, 2022; Lee et al., 2023) extended these ideas and proposed to learn the forward process. However, although NDMs and these works are similar in spirit, they differ in that they optimize the forward process specifically to obtain straight generative trajectories, while in our approach we optimize learnable forward process to minimize variational bound on NLL, which not necessarily leads to straight generative trajectories.
@@ A.3 Reverse SDE and ODE
In the case where $\mathbb{\nu}_{t}$ is a constant, the dynamics become deterministic and can be described by ordinary differential equations (ODEs). In our experiments, we utilize these ODEs to model the generative process as a continuous normalizing flow (Chen et al., 2018; Grathwohl et al., 2018) and estimate densities.
@@ Appendix B Connections with other works
DDPM (Ho et al., 2020) / DDIM (Song et al., 2020a) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$
Flow Matching OT (Lipman et al., 2022) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$ $\alpha_{t}=t$, $\sigma_{t}=1-(1-\sigma_{\mathrm{min}})t$
@@
IHDM (Rissanen et al., 2022) $\mathcal{N}\Big{(}\mathbf{z}_{t};Ve^{-\Lambda t}V^{T}\mathbf{x},\sigma^{2}I\Big{)}$ $Ve^{-\Lambda t}V^{T}\mathbf{x}$ $\alpha_{t}=1$, $\sigma_{t}=\sigma$, $\sigma$ is fixed
Blurring Diffusion (Hoogeboom & Salimans, 2022) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}e^{-\Lambda t}V^{T}\mathbf{x},\sigma_{t}^{2}I\Big{)}$ $e^{-\Lambda t}V^{T}\mathbf{x}$ $p(x|z_{0})=\mathcal{N}\Big{(}x;aVz_{0},\sigma\Big{)}$
@@
f-DM (Gu et al., 2022) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x}_{t},\sigma_{t}^{2}I\Big{)}$ $\mathbf{x}_{t}=\frac{(t-\tau_{k})\hat{x}^{k}+(\tau_{k+1}-t)x^{k}}{\tau_{k+1}-\tau_{k}}$, where $\tau_{k}\leq t<\tau_{k+1}$ $x^{k}&=f_{0:k}(x)\\ \hat{x}^{k}&=\begin{cases}g_{k}(f_{k+1}(x^{k})),&\text{if $k<K$},\\ x^{k},&\text{if $k=K$}.\end{cases}\end{aligned}$
@@ B.1 Diffusion in latent space
The concept of a learnable forward process is not entirely new. In some sense models that run a diffusion process in the latent space of a VAE (Vahdat et al., 2021; Rombach et al., 2022), a hierarchical VAE (Gu et al., 2022), or a Flow model (Kim et al., 2022) can be viewed as diffusion models with a learnable forward process. These models optimize the mapping to the latent space. Consequently, projecting the diffusion generative dynamic from the latent to the data space introduces a novel, nonlinear, and learnable generative dynamic. However, these models still rely on conventional diffusion in the latent space.
@@ B.3 Stochastic Interpolants
Secondly, in NDM, the model is trained by optimizing the variational bound on the likelihood, while Stochastic Interpolants are trained by optimizing the generalization of the Flow Matching objective (Lipman et al., 2022).

### v3

@@ 1 Introduction
Generative models are a powerful class of probabilistic machine learning models with a wide range of applications from e.g. art and music to medicine and physics (Tomczak, 2022; Creswell et al., 2018; Papamakarios et al., 2021; Yang et al., 2022). Generative models learn to mimic the underlying probability distribution of a given data set and can generate novel samples that are similar to the original data. They can for example be used for data augmentation, as well as for unsupervised learning.
Diffusion models have emerged as a family of generative models that excel at several generative tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020). They parameterize the data model through an iterative refinement process, the reverse process, that builds up the data step-by-step from pure noise. For training purposes an auxiliary noising process, the forward process, is introduced that successively adds noise to data. The reverse process is then optimized to resemble the forward process. Despite success in various domains (Sohl-Dickstein et al., 2015; Ho et al., 2020; Saharia et al., 2021; Popov et al., 2021; Watson et al., 2022; Trippe et al., 2023), a key limitation of most existing diffusion models is that they rely on a fixed and pre-specified forward process that is unable to adapt to the specific task or data at hand. At the same time there are many works (Hoogeboom & Salimans, 2023; Rombach et al., 2022; Lipman et al., 2023) that improve performance of diffusion models by modifications of the forward processes.
@@
The main motivation for NDMs is improved likelihood and density estimation, crucial for applications to compression (MacKay, 2003), semi-supervised learning (Dai et al., 2017), adversarial purification (Song et al., 2017), and many others. However, for completeness we also study the impact of NDMs on image generation quality. We find that for small to medium number of steps NDMs achieves better image generation quality than denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020), and comparable results for a large number of steps.
@@
DDPM (Ho et al., 2020) / DDIM (Song et al., 2021a) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$
Flow Matching OT (Lipman et al., 2023) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$ $\alpha_{t}=t$, $\sigma_{t}=1-(1-\sigma_{\mathrm{min}})t$
@@ 2 Background
The combination of the forward process $q$ and the reverse process $p_{\theta}$ is a form of (hierarchical) variational autoencoder (Kingma & Welling, 2014; Rezende et al., 2014). Therefore, it can be trained by optimizing the usual variational bound on the negative log-likelihood. In the case of diffusion models, it can be written as follows (see Section A of Ho et al. (2020)):
@@
By choosing a specific value for $\tilde{\sigma}_{s|t}^{2}$, we can obtain equality between the processes of DDPM and DDIM (see section 4.1 of Song et al. (2021a)). Furthermore, as Song et al. (2021c) demonstrated, when the number of steps $T$ in DDPM goes to infinity, we can transition to continuous time. In this scenario, the reverse process can be described using a Stochastic Differential Equation (SDE):
@@ 3 Neural diffusion models
In theory, we can view diffusion models as a special type of hierarchical variational autoencoders (VAE). From this perspective, the conventional diffusion model resembles a VAE with a fixed variational distribution, in which the latent variables are inferred using scaling of data points and injecting of Gaussian noise. Such a formulation limits diffusion models in terms of the flexibility of the latent space. Introducing a more flexible (and learnable) distribution of latent variables could effectively reduce the gap between the log-likelihood and the variational bound. In practical terms, a more flexible forward process might simplify the task of learning the reverse (generative) process, thereby enhancing model quality. To overcome this limitation of conventional diffusion models, we propose a general form of transformations of data that allows to define and learn distributions on the latent space.
@@ 3.1 Model definition and variational objective
for $0\leq s\leq t\leq T$ where $\tilde{\sigma}_{s|t}^{2}\leq\sigma_{s}^{2}$ is a design choice. Using this posterior we can define an implicit forward process according to (1) (see Figure 1(b)). This forward process provides access to both marginal and posterior distributions just like in the DDIM framework (Song et al., 2021a). The corresponding NDM variational objective has the following form:
@@ 3.2 Continuous time NDMs
Similar to training a discrete time NDM, we can train a continuous time NDMs by sampling time. In our experiments we use importance sampling (Song et al., 2021b) and sample time from a distribution proportional to $\frac{1}{g^{2}(t)}$.
@@ 4 Experiments
Table 2: We report Negative Log-Likelihood (NLL) in Bits Per Dimension (BPD) on the test sets for CIFAR10, ImageNet 32x32, and ImageNet 64x64. Our results were obtained using the continuous-time formulation of our model, integrated via the corresponding Ordinary Differential Equation (ODE), as detailed in Section 3.2.
@@
Score SDE (Song et al., 2021c) $2.99$
Score Flow (Song et al., 2021b) $2.83$ $3.76$
@@ 4.1 Implementation details
We demonstrate NDMs with learnable transformations on the MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) and CelebA-HQ-256 (Karras et al., 2017) datasets. In all experiments we use same neural network architectures to parameterize both the generative process and the transformations $F_{\varphi}$. In experiments with images we use the U-Net architecture from Dhariwal & Nichol (2021). To ensure consistency with Song et al. (2021c, b), we apply horizontal flipping as a data augmentation technique for training models on CIFAR-10 and ImageNet. Unless otherwise stated, we utilize the DDPM variance-preserving schedule of noise injection for $\alpha_{t}$ and $\sigma^{2}_{t}$. For density estimation of discrete data we use uniform dequantization.
In the experiments we report negative log-likelihood (NLL) in bits per dimension (BPD), negative evidence lower bound (NELBO) (8), and sample quality as measured by the Frechet Inception Distance (FID) (Heusel et al., 2017). We calculate NLL by integrating the corresponding ODEs using the RK45 solver from Dormand & Prince (1980), and both NLL and NELBO are calculated on test data. For FID we report the score computed using 50k generated images.
@@ 4.2 Learned transformations
Let us examine some of the transformations that NDM learns. Figure 2(a)-2(c) illustrates the transformations that NDM learns for the 2D checkerboard distribution, MNIST, and CIFAR-10 datasets. For the checkerboard, we observe that $F_{\varphi}$ learns to transform the interleaved pattern into a non-interleaved one. In the case of the grayscale digits of the MNIST dataset, $F_{\varphi}$ learns to highlight the distinctive features of the numbers. It thickens the lines and even creates bubbles at the corners. For the color images of CIFAR-10, $F_{\varphi}$ learns to increase the image contrast. In all cases, our model learns a way to simplify the data distribution. These transformations may enable the reverse process to transition more smoothly from simple distributions to complex ones.
@@ 4.3 Image generation
See Appendix D for additional results and ablation studies.
@@ 5 Related work
NDMs build on diffusion probabilistic models originally proposed by Sohl-Dickstein et al. (2015), which can be considered as an instance of (hierarchical) variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014). Recently, the theory of diffusion models was extended to deterministic sampling (Song et al., 2021a) and continuous time (Song et al., 2021c). These results allowed to reach impressive performance in image generation tasks (Ho et al., 2020; Song et al., 2021c; Dhariwal & Nichol, 2021; Kingma et al., 2021). However, most existing diffusion models have a significant limitation in that they rely on a pre-specified and simple noise injection process that is unable to adapt to the specific task or data at hand. To overcome this, researchers have explored ways to generalize diffusion models.
Various papers have since proposed ways to speed up sampling from diffusion models. Tachibana et al. (2021) and Liu et al. (2022) proposed alternative SDE and ODE solvers. Xiao et al. (2021) proposed replacing simple Gaussian distributions at each generation step with distributions learned by GANs (Goodfellow et al., 2014). Some works proposed methods like iterative distillation with a reduction in the number of steps (Salimans & Ho, 2022) and iterative straightening of trajectories (Liu et al., 2023; Liu, 2022). While these methods change the generative process, they are compatible with NDMs.
Several papers proposed constructing the process of data corruption not by noise injection, but rather by blurring (Rissanen et al., 2023; Daras et al., 2022; Hoogeboom & Salimans, 2023) or through another linear transformation (Singhal et al., 2023). Another line of work modifies directly the dynamics of diffusion models through mapping the data into the latent space of VAE (Vahdat et al., 2021; Rombach et al., 2022), hierarchical VAE (Gu et al., 2023) or normalizing flow (Kim et al., 2022) models and then runs standard linear diffusion. As we demonstrate in Tables 1, these arise as distinct special cases of NDMs for specific choices of the transformation $F_{\varphi}$.
@@
Inspired by diffusion models, several works (Lipman et al., 2023; Neklyudov et al., 2022) have proposed simulation-free objectives for training continuous normalizing flows. These approaches are similar to diffusion models as they rely on the idea of reversing a predefined corruption process. Later, some works (Albergo & Vanden-Eijnden, 2023; Lee et al., 2023) extended these ideas and proposed to learn the forward process. However, although NDMs and these works are similar in spirit, they differ in that they optimize the forward process specifically to obtain straight generative trajectories, while in our approach we optimize learnable forward process to minimize variational bound on NLL, which not necessarily leads to straight generative trajectories.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3 Reverse SDE and ODE
In the case where $\mathbb{\nu}_{t}$ is a constant, the dynamics become deterministic and can be described by ordinary differential equations (ODEs). In our experiments, we utilize these ODEs to model the generative process as a continuous normalizing flow (Chen et al., 2018; Grathwohl et al., 2019) and estimate densities.
@@ Appendix B Connections with other works
DDPM (Ho et al., 2020) / DDIM (Song et al., 2021a) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$
Flow Matching OT (Lipman et al., 2023) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x},\sigma^{2}_{t}I\Big{)}$ $\mathbf{x}$ $\alpha_{t}=t$, $\sigma_{t}=1-(1-\sigma_{\mathrm{min}})t$
@@
IHDM (Rissanen et al., 2023) $\mathcal{N}\Big{(}\mathbf{z}_{t};Ve^{-\Lambda t}V^{T}\mathbf{x},\sigma^{2}I\Big{)}$ $Ve^{-\Lambda t}V^{T}\mathbf{x}$ $\alpha_{t}=1$, $\sigma_{t}=\sigma$, $\sigma$ is fixed
Blurring Diffusion (Hoogeboom & Salimans, 2023) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}e^{-\Lambda t}V^{T}\mathbf{x},\sigma_{t}^{2}I\Big{)}$ $e^{-\Lambda t}V^{T}\mathbf{x}$ $p(x|z_{0})=\mathcal{N}\Big{(}x;aVz_{0},\sigma\Big{)}$
@@
f-DM (Gu et al., 2023) $\mathcal{N}\Big{(}\mathbf{z}_{t};\alpha_{t}\mathbf{x}_{t},\sigma_{t}^{2}I\Big{)}$ $\mathbf{x}_{t}=\frac{(t-\tau_{k})\hat{x}^{k}+(\tau_{k+1}-t)x^{k}}{\tau_{k+1}-\tau_{k}}$, where $\tau_{k}\leq t<\tau_{k+1}$ $x^{k}&=f_{0:k}(x)\\ \hat{x}^{k}&=\begin{cases}g_{k}(f_{k+1}(x^{k})),&\text{if $k<K$},\\ x^{k},&\text{if $k=K$}.\end{cases}\end{aligned}$
@@ B.1 Diffusion in latent space
The concept of a learnable forward process is not entirely new. In some sense models that run a diffusion process in the latent space of a VAE (Vahdat et al., 2021; Rombach et al., 2022), a hierarchical VAE (Gu et al., 2023), or a Flow model (Kim et al., 2022) can be viewed as diffusion models with a learnable forward process. These models optimize the mapping to the latent space. Consequently, projecting the diffusion generative dynamic from the latent to the data space introduces a novel, nonlinear, and learnable generative dynamic. However, these models still rely on conventional diffusion in the latent space.
@@ B.3 Stochastic Interpolants
Secondly, in NDM, the model is trained by optimizing the variational bound on the likelihood, while Stochastic Interpolants are trained by optimizing the generalization of the Flow Matching objective (Lipman et al., 2023).